OpenStudy (anonymous):

What is the distance from (−2, 4) to (0, −6)? Round your answer to the nearest hundredth.

Mehek (mehek14):

\(D = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\)

Mehek (mehek14):

so 0 -(-2) = 2 \(2^2=4\) -6 - 4 = -10 \(-10^2 = 100\) \(\sqrt{104} = 10.20\)

OpenStudy (anonymous):

thank you so much